Write a rule for a function, P

P, that gives the perimeter of a square in inches when the side length is x

x inches.


1
Expert's answer
2022-02-04T14:15:42-0500

Let x be the length (in inches) of the side of a square, then the perimeter "\displaystyle
P(x)" of the square (in inches) is;

"\displaystyle
P(x)=x+x+x+x=4x"